OpenStudy (anonymous):

Sophia invested some money in a bank at a fixed rate of interest compounded annually. The equation below shows the value of her investment after x years. f(x) = 500(1.05)x What was the average rate of change of the value of Sophia's investment from the second year to the fourth year? 14.13 dollars per year 28.25 dollars per year 50.00 dollars per year 56.50 dollars per year

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

i think the first thing to do is calculate for the second year and for the fourth year

OpenStudy (wolf1728):

I think it is important to get the formula correct also: instead of this f(x) = 500(1.05) Use this total = principal * (1 + rate) ^time (notice we need exponentiation here.)

OpenStudy (wolf1728):

Well, since everyone has left, I might as well post the answer. total = principal * (1 + rate) ^time For year zero this calculates to 500.00 = 500.00 * (1.05)^0 525.00 = 500.00 * (1.05)^1 551.25 = 500.00 * (1.05)^2 578.81 = 500.00 * (1.05)^3 607.75 = 500.00 * (1.05)^4 Year 4 - Year 2 amount = 607.75 -551.25 which equals 56.50 per TWO years or (56.50 / 2) = 28.25 per year.

OpenStudy (horrorfreakk):

Thank you, wolf1728!